The perimeter of a rectangle is 22m. The lenght is 5 more than twice the width. Find the dimesions.
L =
W =

Answer by oscargut(686)

L=W+5
and perimeter=p=2(L+W)=22
then L+W=11 and then (W+5)+W=11 then 2W+5=11 then 2W=6 then W=3 and L=11-W=8
Answer: L=8 and W=3